You want to form a value–weighted technology stock index using Apple, Google, and

Intel.  Apple’s adjusted closing price for 2019 is $293.65 and for 2018 is $155.41;

Google’s adjusted closing price for 2019 is $1,337.02 and for 2018 is $1,035.61; Intel’s

adjusted closing price for 2019 is $59.85 and for 2018 is $45.79.

The additional information that you gathered is their market values at the end of 2018 or

beginning of 2019; in billions of  dollars, Apple's market value is $740.93 B, Google's market

value is $790.06 B, and Intel's market value is $198.35 B. What is the annual return for

your value–weighted technology stock index for 2019?
Answer #1

Solution:

MV Weighted technology index
Apple 740.93 740.93 / 1729.34 = 42.84%
Google 790.06 790.93 / 1729.34 = 45.69%
Netflix 198.35 198.35 / 1729.34 = 11.47%
1729.34